The firm’s total cost function is given a follow;

TC = q 2 + wq ; where w =10.


If there are 1000 firms in an industry,

 Find the Short Run Supply Curve of a Firm and industry.

 Also Calculate the Equilibrium Price and Quantity if,


 Qd = 9000 – 200P


 Find out the supply by each firm.


1
Expert's answer
2021-02-10T15:21:34-0500

The short-run individual supply curve is the individual's marginal cost at all points greater than the minimum average variable cost. So, the equation is MC=TC'=2q+10; AVC=(q^2+10q)/q=q+10. 2q+10>q+10 for all q>0.So, short-run individual supply curve is P=2q+10, or q=0.5P-5.

Short-run supply curve of industry is Q=1000q=500P-5000.

Equilibrium price and quantity can be found from equation Qd =Qs; 9000 – 200P=500P-5000; P=14000/700=20. Q=9000-200*20=5000.

The supply by each firm is 5000/1000=5